OpenStudy (anonymous):

Solve: 4(x + 1) + 8 = 24 3.75 −2 5 3

OpenStudy (diyadiya):

4(x+1)+8=24 Open the brackets 4x+4 +8=24 4x+12=24 Subtract 12 from both sides 4x=24-12 4x=12 Divide both sides by 4 4x/4=12/4 x = ?

OpenStudy (anonymous):

4X+4+8=24 4X=12 X=3

OpenStudy (anonymous):

12/4 =3

OpenStudy (diyadiya):

Right ! :)